please help.
The length of one of the equal legs of an isosceles triangle is 8cm less than 4 times the length of the base. If the pereminter is 29cm, find the length of one of the equal legs.

Answer by Earlsdon(6103)

The length of one of the equal sides can be written as:
4x-8 where x is the length of the base.
The perimeter is equal to the sum of the lengths of the three legs of the triangle...so:
(4x-8)+(4x-8)+x = 29cm Simplify and solve for x.
9x-16 = 29 Add 16 to both sides.
9x = 45 Divide both sides by 9
x = 5 cm This is the length of the base.
4x-8 = 4(5)-8 = 20-8 = 12 cm This is the length of one of the equal legs.